What is the molar mass of ammonium carbonate (NH4)2CO3?

A

48.06g/mol

B

96.11g/mol

C

82.09g/mol

D

78.05g/mol

E

192.2g/mol

Grade:12th pass

1 Answers

Mayank Ranka
3 years ago

(NH4)2CO  is the chemical formula of ammonium carbonate, a chemical compound with a molar mass of 96.09 grams per mole

N2=14×(2)=28
H4=1×(4)×2=8
C=12×1=12
O=16×(3)=48
molar mass =28+8+12+48=96